# Optical control of spin-splitting in an altermagnet

### v1

@@ Abstract
Altermagnetism, which has emerged as a new class of magnetism besides ferromagnetism and antiferromagnetism, has drawn a lot of attention in condensed matter physics Šmejkal et al. (2022a, 2020, b). Altermagnets exhibit properties of both ferromagnets (FMs) and antiferromagnets (AFMs). Similarly to AFMs, they have zero net magnetization with local magnetic moments alternating in real space. However, like FMs, they break Kramers’s degeneracy in reciprocal space in the absence of spin-orbit coupling (SOC). Instead of spatial translation and spin-inversion, the opposite spin sublattices in an altermagnet are connected by spatial rotation. Due to the breaking of time-reversal symmetry in momentum space, there are several unconventional anomalous magnetic responses predicted in altermagnets, such as the anomalous Hall effect Šmejkal et al. (2020); Feng et al. (2022); Gonzalez Betancourt et al. (2023) and the magneto-optical Kerr effect Gray et al. (2024), which have been verified by experiments.
The spin-splitting in altermagnets originates from anisotropic exchange interactions. Many materials with the potential to host altermagnetism have been identified Šmejkal et al. (2022a); Guo et al. (2023). However, the predicted spin-splitting of most of them is less than 0.50 eV. Few exceptions are RuO2, CrSb and MnTe. The maximum spin-splitting of 1.4 eV is predicted for RuO2, followed by CrSb and MnTe with spin-splittings of 1.2 eV and 1.1 eV, respectively Šmejkal et al. (2022a).
After the discovery of alternative magnetism, most research efforts have focused on identifying promising materials. Although spin-splitting plays a pivotal role in most of the predicted applications of altermagnet, including spintronics, the possible ways to manipulate and control the spin-splitting gap in these materials remain largely unexplored. Advances in ultrafast science have resulted in promising routes to manipulate and control the properties of materials de la Torre et al. (2021); Rajpurohit et al. (2022, 2020). Optical tuning of spin-splitting gaps can offer exciting opportunities to modify the altermagnetic properties on ultrafast timescales.
@@
Figure 1: Altermagnetic ground state of the TB model described in the text before photoexcitation. (a) Spin and orbital order hosting altermagnetism: The rectangles describe the spin and orbital polarization of the TM sites. Blue and red specify the two opposite spin orientations directions. The orientation, horizontal vs. vertical, encodes the orbital polarization of the $e_{g}$-electron in the xy-plane along the x and y directions, respectively. (b) Average local magnetic moment $\langle|\mu_{R}|\rangle$ (bottom) and octahedral modes $\langle|Q_{2,R}|\rangle$ (top) in the $U/J-U/t$ plane. (c-d) Density of states (DOS) (c) and band structure (d) projected on local orbitals $|\Theta_{l}\rangle$ (green) and $|\Theta_{u}\rangle$ (orange). The left and right DOS shows spin-up and spin-down contributions. The thickness of the color represents the intensity of each component in the band structure. (e) Band structure of the altermagnetic state projected onto spin-up (blue) and spin-down (red) components, with the thickness of the colors representing the intensity of each component. The band structures are plotted along high symmetry points $M(\pi,\pi)-Y(0,\pi)-\Gamma(0,0)-X(\pi,0)-M(\pi,\pi)$. The plots in (b) use $U{=}2.50$ eV. The plots c-e use $t=0.833$ eV, $U/t=3.0$ and $U/J=7.0$.
@@
is conveniently expressed in terms of the on-site terms of the one-particle-reduced density matrix ${\hat{\rho}}$ with the matrix elements defined as
@@
Phononic subsystem: The phononic subsystem consists of two octahedral modes of oxygen around every TM-site $R$: a Jahn-Teller (JT) mode $Q_{2,R}$ and a breathing mode $Q_{1,R}$ Sotoudeh et al. (2017). These phonon modes are defined by the displacement vectors of oxygen ions from their equilibrium positions. The JT mode is an asymmetric expansion of an octahedron in the plane, i.e. expansion along $x$ and contraction along the $y$-direction. The breathing mode is the isotropic expansion of the oxygen octahedron.
@@
Electron-phonon coupling: We consider strong coupling between the eg-electron at TM-sites to the local modes $Q_{1,R}$ and $Q_{2,R}$ Sotoudeh et al. (2017) (see SI). In 3d TM oxides like manganites and nickelates, this type of el-ph interaction is another mechanism besides the Coulomb interaction, which causes long-range orbital ordering Yunoki et al. (2000); Şen et al. (2010); Rajpurohit et al. (2020). The el-ph coupling $E_{el-ph}$ in the model is defined as
@@
We take model parameters from our previous study as a reference Sotoudeh et al. (2017); Rajpurohit et al. (2020, 2020). The reference model parameters are $g_{br}=2.988$ eV/Å, $K_{br}=10.346$ eV/Å2, $g_{JT}=2.113$ eV/Å, $K_{JT}=5.173$ eV/Å2, $U/t=4.29$, $U/J=2.63$ and $t_{hop}=t=0.585$ eV. In the present study, we are interested in the effect of electronic interactions and correlations on the altermagnetic properties, we vary $U$, $J$ and $t_{hop}$ while keeping the other parameters fixed.
Altermagnetic ground-state: We investigate the region of phase diagram with the altermagnetic ground-state, shown in Figure 1-a. We consider one electron per TM-site. Figures 1-b displays the changes in the average local mode $\langle Q_{1,R}\rangle$ and the magnetic moment $\langle|m_{R}|\rangle$ at the TM-sites in the altermagnetic ground-state as a function of $U/t$ and $U/J$ for $U{=}2.50$ eV. For $U/J{<}2.5$ and $U/t{<}2.0$, the system exhibits a non-magnetic metallic phase characterized by TM-sites with zero magnetic and orbital moment. As $U/J$ and $U/t$ increase, the system undergoes metallic to an insulating altermagnetic phase transition. In the altermagnetic phase, the TM-sites have finite magnetic moments and eg-orbital polarization. These local magnetic moments are arranged in antiferromagnetic (AFM) order. The local eg-orbital polarization forms long range staggered orbital ordering as shown in Figure 1-a. This altermagnetic state remains stable for $6.02.5{<}U/J{<}$ 6.0 and $6.02.0{<}U/t{<}$ 6.0.
@@
Figure 2: Spin-splitting gaps in the altermagnetic ground state as functions of $U/t$ (a) and $U/J$ (b). The color codes are the same as 1-e. The other model parameters are set to reference values, which are specified in the main text. The vertical dashed line in (b) indicates $U/J{=}7.0$ used to study photoexcitation.
@@
Photo-excitation study with rt-TDDFT: To study the evolution of spin-splitting under photo-excitation in the altermagnetic state described above, we employ a simulation framework similar to real-time time-dependent density-functional theory (rt-TDDFT) formalism Runge and Gross (1984) based on the model defined in Eq.1. The one-particle wavefunctions of eg-electrons are propagated according to the time-dependent Schrödinger equation. For the sake of simplicity, the atoms are kept frozen while studying the photoexcitation. The photoexcitation is described by an explicit time-dependent vector potentials $\vec{A}(t)=\vec{e}_{s}\omega\rm{Im}(A_{o}e^{-i\omega t})g(t)$ which is coupled to the electrons using Peierls substitution method Peierls (1933). Here, $A_{o}$ is the amplitude of the vector potential and $\omega$ is angular frequency. The envelope $g(t)$ of the light pulse is a Gaussian with a width of 30 fs. The polarization $\vec{e}_{s}$ of light is along the $\hat{x}+\hat{y}$ or the $\hat{x}-\hat{y}$ direction. The unit cell consists of $N{=}4$ TM-sites and 8 oxygen atoms. The simulations are performed with periodic boundary conditions on a 24x24 k-point grid.
@@
Figure 3-b shows the relative change in spin-splitting $\Delta^{s}(t_{f})$$/\Delta^{s}(t{=}0)$ at the high-symmetry point X at $t_{f}{=}50$ fs as a function of photon energy $\hbar\omega$. The largest changes $\Delta^{s}(t_{f})/\Delta^{s}(t{=}0)$ of the spin-splitting gaps occur for photon energies near the bottom of the broad absorption peak, namely at $h\omega_{p}$=2.39 eV.
@@
Let us rationalize the findings described above on the level of local orbitals. For the sake of the argument, we consider here only the diagonal elements of the one-particle reduced density matrix. The photo excitation with photon energy at the absorption maximum transfers electrons from the lower, majority spin orbital of one site to the upper, minority spin orbital of a neighboring site. This implies that, on each site, the occupation change $\delta n_{l,\downarrow,R}$ of the filled orbital equals $-N^{exci}$, while the occupation $\delta n_{u,\uparrow,R}$ of the upper, minority-spin orbital changes by $+N^{exci}$. Figure 4-a shows the electron distribution in the photo-excited state at time $t{=}75$ fs for $N^{exci}=20.30\%$. The excited electrons predominantly occupy the upper conduction band on the $X-M$ and the $Y-M$ high-symmetry lines in reciprocal space.
The orbital energies are obtained using Janak’s theoremJanak (1978), as the derivative of total energy with respect to occupations $n_{\sigma\alpha}$ (see SI for more information). Hence, the occupation changes $\delta n_{\alpha,\sigma,R}$ result in a response $\delta\bar{\epsilon}_{\beta,\sigma^{\prime},R}$ of the orbital energies of the form $\delta\bar{\epsilon}_{\alpha,\sigma,R}=\sum_{\beta,\sigma^{\prime}}\frac{\partial^{2}E_{coul}}{\partial n_{\alpha,\sigma,R}\partial n_{\beta,\sigma^{\prime},R}}\delta{n}_{\beta,\sigma^{\prime},R}$.
@@
Our study demonstrates that spin-splittings of altermagnets can be modified and controlled by optical excitations. These changes occur on ultrafast time scales. Promising candidates to realize this effect are 3d transition-metal compounds. The temporal change in spins-splittings predicted here can experimentally be probed by time-resolved spin angle-resolved photoemission spectroscopy (SARPES) Boschini et al. (2024); Sobota et al. (2021); Rohwer et al. (2011).
@@
Acknowledgments: Theory and simulation were supported by the Computational Materials Sciences (CMS) Program funded by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. Data analysis was provided by the User Program of the Molecular Foundry, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. S.R and T.O. are supported by the Computational Materials Sciences Program funded by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. This work is funded in parts by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) 217133147/SFB1073, projects B03 and C03.

### v2

@@ Abstract
Altermagnetism, which has emerged as a new class of magnetism besides ferromagnetism and antiferromagnetism, has drawn a lot of attention in condensed matter physics Smejkal2022; Smejkal2020; Smejkal2022_2. Altermagnets exhibit properties of both ferromagnets (FMs) and antiferromagnets (AFMs). Similarly to AFMs, they have zero net magnetization with local magnetic moments alternating in real space. However, like FMs, they break Kramers’s degeneracy in reciprocal space in the absence of spin-orbit coupling (SOC). Instead of spatial translation and spin-inversion, the opposite spin sublattices in an altermagnet are connected by spatial rotation. Due to the breaking of time-reversal symmetry in momentum space, there are several unconventional anomalous magnetic responses predicted in altermagnets, such as the anomalous Hall effect Smejkal2020; Feng2022; Betancourt2023 and the magneto-optical Kerr effect Isaiah2024, which have been verified by experiments.
The spin-splitting in altermagnets originates from anisotropic exchange interactions. Many materials with the potential to host altermagnetism have been identified Smejkal2022; Guo2023. However, the predicted spin-splitting of most of them is less than 0.50 eV. Few exceptions are RuO2, CrSb and MnTe. The maximum spin-splitting of 1.4 eV is predicted for RuO2, followed by CrSb and MnTe with spin-splittings of 1.2 eV and 1.1 eV, respectively Smejkal2022.
After the discovery of alternative magnetism, most research efforts have focused on identifying promising materials. Although spin-splitting plays a pivotal role in most of the predicted applications of altermagnet, including spintronics, the possible ways to manipulate and control the spin-splitting gap in these materials remain largely unexplored. Advances in ultrafast science have resulted in promising routes to manipulate and control the properties of materials Torre2021; Rajpurohit2022; Rajpurohit2020. Optical tuning of spin-splitting gaps can offer exciting opportunities to modify the altermagnetic properties on ultrafast timescales.
@@
Figure 1: Altermagnetic ground state of the TB model described in the text before photoexcitation. (a) Spin and orbital order hosting altermagnetism: The rectangles describe the spin and orbital polarization of the TM sites. Blue and red specify the two opposite spin orientations directions. The orientation, horizontal vs. vertical, encodes the orbital polarization of the $e_{g}$-electron in the xy-plane along the x and y directions, respectively. (b) Average local magnetic moment $\langle|\mu_{R}|\rangle$ (bottom) and octahedral modes $\langle|Q_{2,R}|\rangle$ (top) in the U/JU/tU/J-U/t plane. (c-d) Density of states (DOS) (c) and band structure (d) projected on local orbitals $|\Theta_{l}\rangle$ (green) and $|\Theta_{u}\rangle$ (orange). The left and right DOS shows spin-up and spin-down contributions. The thickness of the color represents the intensity of each component in the band structure. (e) Band structure of the altermagnetic state projected onto spin-up (blue) and spin-down (red) components, with the thickness of the colors representing the intensity of each component. The band structures are plotted along high symmetry points $M(\pi,\pi)-Y(0,\pi)-\Gamma(0,0)-X(\pi,0)-M(\pi,\pi)$. The plots in (b) use U=2.50U{=}2.50 eV. The plots c-e use $t=0.833$ eV, $U/t=3.0$ and $U/J=7.0$.
@@
is conveniently expressed in terms of the on-site terms of the one-particle-reduced density matrix $^{\hat{\rho}}$ with the matrix elements defined as
@@
Phononic subsystem: The phononic subsystem consists of two octahedral modes of oxygen around every TM-site $R$: a Jahn-Teller (JT) mode $Q_{2,R}$ and a breathing mode $Q_{1,R}$ Sotoudeh2017. These phonon modes are defined by the displacement vectors of oxygen ions from their equilibrium positions. The JT mode is an asymmetric expansion of an octahedron in the plane, i.e. expansion along $x$ and contraction along the $y$-direction. The breathing mode is the isotropic expansion of the oxygen octahedron.
@@
Electron-phonon coupling: We consider strong coupling between the eg-electron at TM-sites to the local modes $Q_{1,R}$ and $Q_{2,R}$ Sotoudeh2017 (see SI). In 3d TM oxides like manganites and nickelates, this type of el-ph interaction is another mechanism besides the Coulomb interaction, which causes long-range orbital ordering Yunoki2000; Sen2010; Rajpurohit2020. The el-ph coupling $E_{el-ph}$ in the model is defined as
@@
We take model parameters from our previous study as a reference Sotoudeh2017; Rajpurohit2020_2; Rajpurohit2020. The reference model parameters are $g_{br}=2.988$ eV/Å, $K_{br}=10.346$ eV/Å2, $g_{JT}=2.113$ eV/Å, $K_{JT}=5.173$ eV/Å2, $U/t=4.29$, $U/J=2.63$ and $t_{hop}=t=0.585$ eV. In the present study, we are interested in the effect of electronic interactions and correlations on the altermagnetic properties, we vary $U$, $J$ and $t_{hop}$ while keeping the other parameters fixed.
Altermagnetic ground-state: We investigate the region of phase diagram with the altermagnetic ground-state, shown in Figure 1-a. We consider one electron per TM-site. Figures 1-b displays the changes in the average local mode $\langle Q_{1,R}\rangle$ and the magnetic moment $\langle|m_{R}|\rangle$ at the TM-sites in the altermagnetic ground-state as a function of $U/t$ and $U/J$ for U=2.50U{=}2.50 eV. For U/J<2.5U/J{<}2.5 and U/t<2.0U/t{<}2.0, the system exhibits a non-magnetic metallic phase characterized by TM-sites with zero magnetic and orbital moment. As $U/J$ and $U/t$ increase, the system undergoes metallic to an insulating altermagnetic phase transition. In the altermagnetic phase, the TM-sites have finite magnetic moments and eg-orbital polarization. These local magnetic moments are arranged in antiferromagnetic (AFM) order. The local eg-orbital polarization forms long range staggered orbital ordering as shown in Figure 1-a. This altermagnetic state remains stable for 2.5<U/J<6.02.5{<}U/J{<}6.0 and 2.0<U/t<6.02.0{<}U/t{<}6.0.
@@
Figure 2: Spin-splitting gaps in the altermagnetic ground state as functions of $U/t$ (a) and $U/J$ (b). The color codes are the same as 1-e. The other model parameters are set to reference values, which are specified in the main text. The vertical dashed line in (b) indicates U/J=7.0U/J{=}7.0 used to study photoexcitation.
@@
Photo-excitation study with rt-TDDFT: To study the evolution of spin-splitting under photo-excitation in the altermagnetic state described above, we employ a simulation framework similar to real-time time-dependent density-functional theory (rt-TDDFT) formalism Runge1984 based on the model defined in Eq.1. The one-particle wavefunctions of eg-electrons are propagated according to the time-dependent Schrödinger equation. For the sake of simplicity, the atoms are kept frozen while studying the photoexcitation. The photoexcitation is described by an explicit time-dependent vector potentials $\vec{A}(t)=\vec{e}_{s}\omega\rm{Im}(A_{o}e^{-i\omega t})g(t)$ which is coupled to the electrons using Peierls substitution method Peierls1933. Here, $A_{o}$ is the amplitude of the vector potential and $\omega$ is angular frequency. The envelope $g(t)$ of the light pulse is a Gaussian with a width of 30 fs. The polarization $\vec{e}_{s}$ of light is along the $\hat{x}+\hat{y}$ or the $\hat{x}-\hat{y}$ direction. The unit cell consists of N=4N{=}4 TM-sites and 8 oxygen atoms. The simulations are performed with periodic boundary conditions on a 24x24 k-point grid.
@@
Figure 3-b shows the relative change in spin-splitting $\Delta^{s}(t_{f})$$\Delta^{s}(t{=}0)$ at the high-symmetry point X at $t_{f}{=}50$ fs as a function of photon energy $\hbar\omega$. The largest changes $\Delta^{s}(t_{f})/\Delta^{s}(t{=}0)$ of the spin-splitting gaps occur for photon energies near the bottom of the broad absorption peak, namely at $h\omega_{p}$=2.39 eV.
@@
Let us rationalize the findings described above on the level of local orbitals. For the sake of the argument, we consider here only the diagonal elements of the one-particle reduced density matrix. The photo excitation with photon energy at the absorption maximum transfers electrons from the lower, majority spin orbital of one site to the upper, minority spin orbital of a neighboring site. This implies that, on each site, the occupation change $\delta n_{l,\downarrow,R}$ of the filled orbital equals $N^{exci}$, while the occupation $\delta n_{u,\uparrow,R}$ of the upper, minority-spin orbital changes by $N^{exci}$. Figure 4-a shows the electron distribution in the photo-excited state at time t=75t{=}75 fs for $N^{exci}=20.30\%$. The excited electrons predominantly occupy the upper conduction band on the XMX-M and the YMY-M high-symmetry lines in reciprocal space.
The orbital energies are obtained using Janak’s theoremJanak1978, as the derivative of total energy with respect to occupations $n_{\sigma\alpha}$ (see SI for more information). Hence, the occupation changes $\delta n_{\alpha,\sigma,R}$ result in a response $\delta\bar{\epsilon}_{\beta,\sigma^{\prime},R}$ of the orbital energies of the form $\delta\bar{\epsilon}_{\alpha,\sigma,R}=\sum_{\beta,\sigma^{\prime}}\frac{\partial^{2}E_{coul}}{\partial n_{\alpha,\sigma,R}\partial n_{\beta,\sigma^{\prime},R}}\delta{n}_{\beta,\sigma^{\prime},R}$.
@@
Our study demonstrates that spin-splittings of altermagnets can be modified and controlled by optical excitations. These changes occur on ultrafast time scales. Promising candidates to realize this effect are 3d transition-metal compounds. The temporal change in spins-splittings predicted here can experimentally be probed by time-resolved spin angle-resolved photoemission spectroscopy (SARPES) Boschini2024; Sobota2021; Rohwer2011.
@@
Acknowledgments: Theory and simulation were supported by the Computational Materials Sciences (CMS) Program funded by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. Data analysis was provided by the User Program of the Molecular Foundry, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. S.R and T.O. are supported by the Computational Materials Sciences Program funded by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. S.R. was partially supported by funds from the UC National Laboratory Fees Research Program of the University of California, Grant Number L25CR9003. This work is funded in parts by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) 217133147/SFB1073, projects B03 and C03.